your parents are planning to retire in 18 years. They currently have $250,000 and they would like to have $1,000,000 when they retire. What annual rate of interest would they have to earn on their $250,000 in order to reach?
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

A(t) = P(1+rt)
1,000,000 = 250,000(1+r*18)
-----
4 = 1 + 18r
-----
18r = 3
r = 1/6
r = 16.67% or 16 2/3 %
-----------------------
Cheers,
Stan H.

Answer by solver91311(24713)



Just use the rule of 72 for this one. 72 divided by the number of years the money is invested is the interest rate required for the money to double in that number of years. In order for 250K to become 1M, it has to double twice in 18 years, so every 9 years. 72 divided by 9 is 8. So 8% annual interest -- close enough anyway.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism